I've spent a quarter of my life as a boy, a sixth of my life as a young man, half my life as a middle age man; all these years I spent in my native country. Now you must add the past six years I've spent as a new immigrant. How old is he?
X=how old he is
1/4 of x = time as a boy
1/6 of x = young man
1/2 of x = middle age
I'm not sure want to do next. Can you help me please?
My email address is briggs4u@sbcglobal.net

Answer by Alan3354(69443)

I've spent a quarter of my life as a boy, a sixth of my life as a young man, half my life as a middle age man; all these years I spent in my native country. Now you must add the past six years I've spent as a new immigrant. How old is he?
X=how old he is
1/4 of x = time as a boy
1/6 of x = young man
1/2 of x = middle age
-------------------------
x = x/4 + x/6 + x/2 + 6
x = (3x + 2x + 6x)/12 + 6
12x = 11x + 72
x = 72